OpenStudy (anonymous):

Factor this expression. Check by multiplying factors. Medal rewarded for step by step procedures. 2ab-6ac+3b-9c

OpenStudy (cwrw238):

first rearrange 2ab + 3b - 6ac - 9c now can you factor the first 2 and the last 2 terms to get a common factor?

OpenStudy (anonymous):

I don't know how to factor.

OpenStudy (anonymous):

2(a+3)(b-3c)?

OpenStudy (cwrw238):

first factor 2ab + 3b b is common to both terms and 2ab / b = 2a and 3b / b = 3 so we can write 2ab + 3b = b(2a + 3)

OpenStudy (cwrw238):

-6ac - 9c take out -3c and we get -3c(2a + 3)

OpenStudy (anonymous):

Why are we taking out -3c?

OpenStudy (cwrw238):

because its the highest factor which will divide into both -6ac and - 9c

OpenStudy (anonymous):

Alright, I'm following. So, we just put those two terms together?

OpenStudy (cwrw238):

yes so we have b(2a + 3) - 3c(2a + 3) now 2a + 3 is common to both parts so we have (b-3c)(2a+3) and theres your factors

OpenStudy (anonymous):

Okay, I get it, but what do I do when an expression like this comes up? 3ab +3ac+2b^2+2bc

OpenStudy (cwrw238):

take out 3a for the first 2 and 2b for the last 2 terms the highest common factor is called the HCF (or GCF in the States)

OpenStudy (cwrw238):

this ves you 3a(b + c) + 2b(b + c)

OpenStudy (cwrw238):

same priciple as the first question

OpenStudy (anonymous):

So then we have 3a+2b

OpenStudy (anonymous):

correct? @cwrw238

OpenStudy (cwrw238):

thats in one of the brackets the factors are (3a + 2b)(b+ c)

OpenStudy (cwrw238):

if you multiply this out it should come back to the original expression

OpenStudy (anonymous):

Oh yeah, that's right.. I'm forgetting some steps :(

OpenStudy (anonymous):

3a*b = 3ab 2b*c = 2bc 3ab+2bc.. Like that?

OpenStudy (cwrw238):

first multiply the ( b +c) by 3a then by 2b yes

OpenStudy (anonymous):

3ab+3ac+2b^2+2bc

OpenStudy (cwrw238):

3a*b + 3a*c + 2b*b + 2b*c = 3ab+ 3ax + 2b^2 + 2bc

OpenStudy (cwrw238):

yea - 3ab not 3ax - my mistake

OpenStudy (cwrw238):

I mean 3ac - lol!

OpenStudy (anonymous):

Oh well, we got the original expression back so (3a + 2b)(b+ c) is correct?

OpenStudy (anonymous):

Just double checking before I put in.

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

hope this has helped - i gotta go right now

OpenStudy (anonymous):

thanks alot man! I might have you go over some of my work 'cause I'mma go try a few.

OpenStudy (cwrw238):

i'm a ma'am - but thats ok!

OpenStudy (anonymous):

Sorry madam.